OpenStudy (anonymous):

find the discriminant and the determine the nature of the real roots of 3x^2+3x-60=0

OpenStudy (anonymous):

discriminant = b^2 - 4ac = 9 - 4(3)(-60) =9 + 720 =729 since discriminant is greater than zero the roots are real and distinct